OpenStudy (anonymous):

I give medals! A cup of coffee which is initially at a temperature of 159 F is placed in a room which is at a constant temperature of 68 F. In 7 minutes, the coffee has cooled to 141 F. Find the Newton's Law of Cooling formula that models the coffee's temperature in minutes. Keep at least 4 decimal places in your formula for rounded values. Determine to the nearest minute how long it will take the coffee to cool to a temperature of 100 F

OpenStudy (anonymous):

\[T(t)= S+(T _{0}-S)e ^{kt}\] \[141=68+(159-68)e ^{7k}\]

OpenStudy (anonymous):

this is only annoying because you have to work with the differences in the heated temp and the room temp so instead of working with 159 and 141 you work with 91 and 73 then it is just like the last one

OpenStudy (anonymous):

\[\frac{ 73 }{ 91 }=\frac{ 91e ^{7k} }{ 9 }\] \[\frac{ \ln73 }{ 91 }=7k\]

OpenStudy (anonymous):

\[k=\frac{ \ln(73/91) }{ 7}\]

OpenStudy (anonymous):

Just wondering if I'm doing this correct so far

OpenStudy (anonymous):

i assume you mean \[\ln(\frac{73}{91})=7k\] right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

oh yeah, i see from your next post yes, you are right so far

OpenStudy (anonymous):

Mmkay! I just gotta finish it up now

OpenStudy (anonymous):

I got k= -0.0315

OpenStudy (anonymous):

yes, i got that too http://www.wolframalpha.com/input/?i=ln%2873%2F91%29%2F7

OpenStudy (anonymous):

So, would the equation be \[T(t)=68+(T _{0}-68)e ^{-0.0315*t}\]

OpenStudy (anonymous):

\(T_0=159\) so it is \[\large T=68+91e^{-0.0.15t}\]

OpenStudy (anonymous):

\(T_0\) is the initial temperature, not a variable in the formula

OpenStudy (anonymous):

Wow, I'm an idiot! -_- thank you!

OpenStudy (anonymous):

lol yw ok form there? last part is to ste \[\large 91e^{-0.0315t}=32\] and solve for \(t\)

OpenStudy (anonymous):

*set

OpenStudy (anonymous):

Got it! I just have to finish soling it!

OpenStudy (anonymous):

I got t=33

OpenStudy (anonymous):

kk i will too

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

Sweet! I appreciate it! Thanks! :D

OpenStudy (anonymous):

yw again btw you did all the work